OpenStudy (scott208):

If a, b, c, and d are four different digits from 1 to 9, inclusive, then what's the largest possible value for the decimal a.b + c.d? will medaL

OpenStudy (scott208):

@Atsie

Atsie (atsie):

To be very honest, @Scott208 I've not heard of this and neither do I know how to do it. I've read it over and I don't really understand it. I don't want to be misguiding, so I'd trust someone else to be of help to you. I'm really sorry. :/

OpenStudy (scott208):

alright thanks anyways

Atsie (atsie):

You're welcome :) I'll try to get someone else for ya ^.^

imqwerty (imqwerty):

\(a.b+c.d\) \(a+\large\frac{b}{10}\) + \(c +\large\frac{d}{10}\) \(\large \frac{10a+b+10c+d}{10}\) \(\large \frac{a+b+c+d +9a+9c}{10}\) now if we want the maximum value of a.b+c.d then either of a/b/c/d must have values 9/8/7/6 so a+b+c+d is 9+8+7+6=30 so our equation becomes- \(\large \frac{a+b+c+d +9a+9c}{10}\) \(\large \frac{30+9(a+c)}{10}\) \(\large\frac{30}{10}+\large\frac{9(a+c)}{10}\) \(3+\large\frac{9(a+c)}{10}\) now maximum value of a+c can be 9+8=17 so \(3+\large\frac{9(17)}{10}\)

OpenStudy (scott208):

thanks

imqwerty (imqwerty):

np

OpenStudy (scott208):

so 18.3

imqwerty (imqwerty):

yeah

OpenStudy (scott208):

cool

imqwerty (imqwerty):

you can also directly say that for a.b+c.d to be maximum a and c must be the biggest numbers that can be possible so clearly a=9 or 8 and c=8 or 9 similarly d=7 or 6 and b=6 or 7 so the sum would be = 18.3